determine the temperature at which a 17.2 gram sample of ethane gas (C2H6) will occupy a volume of 6.54 liters and have a pressure of 715 mm Hg
1
Expert's answer
2020-05-01T14:41:08-0400

m=17.2 g

M(C2H6)=30.07 g/mol

V=6.54 L

P=715 mm Hg = 0.9408 atm

n=m/M

n=(17.2 g)/(30.07 g/mol)=0.572 mol

Ideal gas law:

PV=nRT

Molar gas constant

R=0.082058 L×atm/(K×mol)

(0.9408 atm)×(6.54 L)=(0.572 mol)×(0.082058 L×atm/(K×mol))×T

T=((0.9408 atm)×(6.54 L))/((0.572 mol)×(0.082058 L×atm/(K×mol)))

T= 131 K

Answer: 131 K